A solution of an unknown nonvolatile nonelectrolyte was prepared by dissolving 0.250 g of the substance in 40.0 g of CCl4. The boiling point of the resultant solution was 0.357oC higher than that of the pure solvent. Calculate the molar mass of the solute.
1
Expert's answer
2020-12-14T14:46:14-0500

Raoult's law states:

"\Delta T = EC_m", where delta T -- difference between boiling points, E -- ebullioscopic constant (5.0 K*kg/mol for CCl4), Cm -- molality of solution.

"C_m = m_x\/Mm_s" , where mx -- mass of the unknown compound, ms -- mass of the solvent.

Combine the two formulas:

"\Delta T= Em_x/Mm_s"

Find the molar mass:

"M = Em_x/m_s\Delta T"

"M = 5*0.25\/(0.04*0.357) = 87.5 (g\/mol)"

Answer: 87.5 g/mol is the molar mass of the solute.